The sum of the age of two sisters is 42. The second sister is three more than twice the first.

Find the age of the OLDER sister. Just give the age - do not include units.

Answer by stanbon(75887)

The sum of the age of two sisters is 42. The second sister is three more than twice the first.
Find the age of the OLDER sister. Just give the age - do not include units.
-----------
Equation:
x + y = 42
y = 2x + 3
------
Substitute for "y" and solve for "x":
x + 2x+3 = 42
3x = 39
x = 13 yrs (younger sister)
Solve for "y":
y = 2x+3 = 2*13+3 = 29 yrs (older sister)
-------------
Cheers,
Stan H.
==============